OpenStudy (anonymous):

solve 3x + 2 = 5x – 6

OpenStudy (kendricklamar2014):

1. Subtract 5x from both sides: 3x + 2 - 5x = 5x - 6 - 5x

OpenStudy (kendricklamar2014):

Which equals: \[-2x + 2 = -6\]

OpenStudy (kendricklamar2014):

2. Subtract 2 from both sides: -2x + 2 - 2 = -6 - 2 Which equals: -2x = -8

OpenStudy (kendricklamar2014):

3. Divide both sides by 2:\[\left(\begin{matrix}-2x \\ -2\end{matrix}\right) = \left(\begin{matrix}-8 \\ -2\end{matrix}\right)\]

OpenStudy (kendricklamar2014):

Which gives you the answer:\[x = 4\]